Find the solution of the given IVP y + 3y + 2y = uz(t); y(0) = 0, y(0) = 1 a. y = et-e-t + uz(t) [] + e-(6+2) +22(6+2) b.
Answer #1

Hexe Criver DE is yll + 3yl + 2y = 42(2), y(0)=0, y/o) = 1 Ult) 1: &<2 Now , taking Laplace transform, we get- L[y] + 3 L[y]O Now [ {f(s)} = 1 • Y14): 1784 +683} tet-e? [ 2 25 3+1 2 (3+2) 4 24 ] = FH + te 2 е 2 Lyfes)} second shifting property. T